# version 0.75

### v4

@@ 6.1.1 Online Gradient Descent
Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). Pseudo-code is given in Algorithm 3.
@@
In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost. This step may result in a point outside of the underlying convex set. In such cases, the algorithm projects the point back to the convex set, i.e. finds its closest point in the convex set. Despite the fact that the next cost function may be completely different than the costs observed thus far, the regret attained by the algorithm is sublinear. This is formalized in the following theorem. Here we define the constants $D$ and $G$ as follows:
@@
The Online Gradient Descent algorithm is straightforward to implement, and updates take linear time given the gradient. However, the projection step may in general take significantly longer via convex optimization.
@@ 6.2 Regret for Control
The most important consideration to explore beyond classical models of control, such as optimal and robust control, is computational. Even in linear control with convex loss functions, the optimal policy can be complex to describe, see exercises section.
A prominent approach to deal with computational difficulty is to consider a less stringent criteria for optimality. Instead of the optimal policy, we can consider all policies in a certain policy class, and try to compete with the optimal policy within the class in terms of regret.
@@
The significance of this definition is that if a simple policy class approximates another more complicated class, then it suffices to consider only the first. The usefulness of this definition stems from its implication for regret minimization, as per definition 1.3, as follows.
@@ 6.3 Expressivity of Control Policy Classes
In this section we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems. The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics.
@@
A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$. Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems. In later chapters of this text we study policies for partially observed systems.
@@ 6.3.1 Linear Policies
The most basic and well-studied policy classes for control of linear dynamical systems are linear policies, i.e. those that choose a control signal as a (fixed) linear function of the state. The most important subcase of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state. This was formalized in Definition 2.4. Formally,
@@
A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator mapping state to control. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
@@ 6.3.2 Disturbance Action Controllers
We next define the most important set of policies in online nonstochastic control, known as the disturbance action controllers, or DAC. The main advantages of DAC as opposed to linear controllers, are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization. Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show below.
@@
Without loss of generality, we take $M_{0}=0$, as we can add a fictitious unit coordinate to the perturbations $\{\mathbf{w}_{t}\}$, and modify the system accordingly. This will allow a constant shift in the control.
@@
The class of DAC always produces a stabilizing control signal whenever $K_{1:T}$ are sequentially stabilizing for the entire system $\{A_{t},B_{t}\}$. This important fact can be seen in the following lemma.
@@
Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below. For simplicity assume $K_{t}=K$ are all equal. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
@@
The above lemma is already a good indication about DAC: they generate controls that at least don’t cause a blow up. But how expressive are these policies in terms of reaching the optimal solution?
One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than linear controllers. Notice that the $K_{t}$ terms for DAC are only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
@@
when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies are executed on the dynamical system $(A,B)$ the iterates guarantee
$\max\{\|\mathbf{x}_{t}^{\pi_{1}}(A,B)-\mathbf{x}_{t}^{\pi_{2}}(A,B)\|,\|\mathbf{u}_{t}^{\pi_{1}}(A,B)-\mathbf{u}_{t}^{\pi_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
@@
$=\sum_{i=0}^{h}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}+Z_{t}$
@@
where $Z_{t}=\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$. The magnitude of this residual term can be bounded, due to the existence of $\frac{\kappa}{\delta}\leq\gamma^{\prime}$ from Exercise 4.6, as
@@
$\leq\sum_{i=h+1}^{t}(1-\delta)^{i+1}\left\|K\mathbf{w}_{t-i}\right\|$
@@
$=\kappa\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
@@
To make a similar observation on the corresponding state sequences, first note that
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
Linear state-feedback policies are a mainstay of control theory both in applications, due to their simplicity, and in theory, owing to their optimality in deterministic and stochastic settings as demonstrated in [Kal60]. The methods for computing such policies are often based in dynamic programming. Often motivated by the need to impose constraints beyond optimality such as robustness, various other policy parametrizations were introduced. An early example is the Youla parametrization, sometimes called Q-parametrization [YJB76]. Later on, the introduction of system level synthesis [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially so given its emphasis on state-space representations. For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20].
The Disturbance Action Control parametrization, introduced in [ABH+19], is similar to system level synthesis. Like SLS, it is a convex parametrization that is amenable to the tools of convex optimization. However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies. This makes it better suited to machine learning.
The study of regret in the context of control has roots going back to the work of [AYS11, AYBK14]. These works, as well as many that followed [CHK+18, AHS19], study regret in a stochastic setting, where the perturbations are stochastic. In such settings, optimal control is well defined. The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19].
@@ 6.5 Exercises
6.1. In this exercise we prove that the optimal solution to a linear control problem can be complex to describe.
@@
    Consider a linear dynamical system in one dimension, and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$. Show that the optimal policy in hindsight can have as many domains as loss functions.
@@
    Consider a similar setting in higher dimensions, and give a bound on the number of domains as a function of the number of loss functions and the dimension.
6.2. Consider the setting of time varying linear dynamical systems, i.e.
@@
Prove that any DAC according to definition 6.5 is stable, i.e. prove the analogue of Lemma 6.6 for time varying systems.
@@
In this chapter, we explore recent advancements in machine learning that have significantly reshaped the control problem. These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with but often outperform traditional approaches. The key characteristics of these modern techniques are as follows:
@@
    Online Learning Over Precomputed Policies Unlike classical control, which relies on precomputing an optimal policy, online control adopts adaptive learning techniques to update policies dynamically. This aligns with the broader framework of adaptive control, where policies evolve in response to observed system behavior.
@@
    Improper Learning via Convex Relaxation A fundamental breakthrough in online nonstochastic control is the use of improper learning, where the algorithm competes against a given policy class while learning from a broader class of policies. By leveraging convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, allowing for efficient algorithms with strong theoretical guarantees.
@@ 7.1 From Optimal and Robust to Online Control
Classical control theory has been primarily shaped by two paradigms: optimal control and robust control. These approaches have been instrumental in various engineering applications, but they rely on specific assumptions about disturbances and system dynamics.
@@
While these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting. Consider the challenge of controlling a drone in an unpredictable environment:
@@
These limitations motivate the need for a more adaptive approach—one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions. This leads us to online nonstochastic control, a fundamentally different paradigm where controllers adapt in real-time without assuming a fixed model for disturbances.
@@ 7.1.1 A Shift to Online Nonstochastic Control
Rather than relying on predefined stochastic or worst-case assumptions, online nonstochastic control introduces a new approach based on sequential decision-making and regret minimization. The core principles are:
@@
To build intuition, consider the problem of controlling an autonomous vehicle in an unknown city. If the vehicle had complete foresight of future traffic conditions, it could compute the optimal route in advance. However, without such knowledge, the vehicle must continuously adjust its route based on real-time observations. Regret measures how far the vehicle’s performance deviates from the best possible path that could have been chosen in hindsight.
@@ 7.2 The Online Nonstochastic Control Problem
In this framework, both the disturbances and cost functions are chosen by an adversary, meaning they may follow no statistical pattern and could even be designed to hinder the controller’s decisions at every step. As a result, the goal is no longer to compute a fixed optimal control policy but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
A fundamental difficulty in control is the computational tractability of finding the best policy. Classical control methods often rely on dynamic programming, which, in general, suffers from the curse of dimensionality. Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems over trajectories. As pointed out by Rockafellar, finding an optimal control policy in arbitrary convex cost settings can be computationally prohibitive, particularly when state and control constraints are involved.
To address this, online nonstochastic control shifts the focus from computing an optimal policy to competing with a benchmark class of policies using regret minimization. Instead of optimizing a predefined objective function, the controller seeks to minimize its total loss relative to the best policy that could have been chosen in hindsight. This approach bypasses the need to solve an intractable dynamic optimization problem upfront while still achieving performance guarantees over time.
@@
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
@@
Improper Learning and Convex Relaxation
@@
This approach is crucial because enforcing exact constraints on the policy space can lead to computational hardness results, as previously discussed. By relaxing the policy representation, for example using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
@@ 7.3 The Gradient Perturbation Controller
The algorithm is described in figure (4), for time varying linear dynamical systems. It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2. The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent. We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
The notation we use to describe the algorithm is as follows. We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time. Note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$. The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
@@
The GPC algorithm comes with a very strong performance guaranteee: it guarantees vanishing regret vs. the best Disturbance Action Control policy belonging to our reference class in hindsight. As per the relationship between DAC and other policies, this implies vanishing regret vs. linear controllers as well. The formal statement is given as follows.
@@
    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable, and have bounded $\ell_{2}$ norm.
@@
As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$. However, to directly apply the OGD algorithm to our setting, we need to ensure that the following conditions hold true:
@@
  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
These facts are formalize in the following two lemmas, that we state here and prove at the end of this section.

### v5

@@ 6.1.1 Online Gradient Descent
Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). The pseudocode is given in Algorithm 3.
@@
In each iteration, the algorithm takes a step from the previous point in the direction of the gradient of the previous cost. This step may result in a point outside the underlying convex set. In such cases, the algorithm projects the point back to the convex set, i.e. finds its closest point in the convex set. Despite the fact that the next cost function may be completely different from the costs observed thus far, the regret attained by the algorithm is sublinear. This is formalized in the following theorem. Here we define the constants $D$ and $G$ as follows:
@@
The Online Gradient Descent algorithm is straightforward to implement, and updates take linear time given the gradient. However, in general, the projection step may take significantly longer via convex optimization.
@@ 6.2 Regret for Control
The most important consideration to explore beyond classical models of control, such as optimal and robust control, is computational. Even in linear control with convex loss functions, the optimal policy can be complex to describe, see the exercises section.
A prominent approach to deal with computational difficulty is to consider a less stringent criterion for optimality. Instead of the optimal policy, we can consider all policies in a certain policy class, and try to compete with the optimal policy within the class in terms of regret.
@@
The significance of this definition is that if a simple policy class approximates another more complicated class, then it suffices to consider only the first. The usefulness of this definition comes from its implications for regret minimization, according to definition 1.3, as follows.
@@ 6.3 Expressivity of Control Policy Classes
In this section, we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems. The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time-varying linear dynamical systems were proposed by Lyapunov as a general methodology to study nonlinear dynamics.
@@
A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$. Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems. In later chapters of this text, we study policies for partially observed systems.
@@ 6.3.1 Linear Policies
The most basic and well-studied policy classes for control of linear dynamical systems are linear policies, i.e. those that choose a control signal as a (fixed) linear function of the state. The most important subcases of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state. This was formalized in Definition 2.4. Formally,
@@
A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator that maps the state to the control. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Frobenius norm by $\|K\|\leq\gamma$.
@@ 6.3.2 Disturbance Action Controllers
We next define the most important set of policies in online nonstochastic control, known as disturbance action controllers, or DAC. The main advantages of DAC, as opposed to linear controllers, are that the cost of control is convex in their parameterization and thus allows for provably efficient optimization. Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show in the following.
@@
Without loss of generality, we take $M_{0}=0$, as we can add a fictitious unit coordinate to the perturbations $\{\mathbf{w}_{t}\}$, and modify the system accordingly. This will allow for a constant shift in the control.
@@
The class of DACs always produces a stabilizing control signal whenever $K_{1:T}$ are sequentially stabilizing for the entire system $\{A_{t},B_{t}\}$. This important fact can be seen in the following lemma.
@@
Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as follows. For simplicity, assume that $K_{t}=K$ are all equal. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
@@
The above lemma is already a good indication about DAC: they generate controls that at least do not cause a blow-up. But how expressive are these policies in terms of reaching the optimal solution?
One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than linear controllers. Notice that the $K_{t}$ terms for the DAC are only assumed to be stabilizing and not necessarily optimal for the actual sequence.
@@
when executed on some dynamical system $(A+BK_{0},B)$, then when the corresponding policies $\pi^{\prime}_{1}$ and $\pi^{\prime}_{2}$, that is, $\pi_{1}$ and $\pi_{2}$ with $K_{0}$ superimposed, are executed on the dynamical system $(A,B)$ the iterates guarantee
[…]
$\|\mathbf{u}_{t}^{\pi^{\prime}_{1}}(A,B)-\mathbf{u}_{t}^{\pi^{\prime}_{2}}(A,B)\|\}\leq(1+\|K_{0}\|)\varepsilon.$
@@
Now, we define $Z_{t}=\sum_{i=h+1}^{t}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}$, and continue as shown.
$\mathbf{u}_{t+1}$ $=\sum_{i=0}^{h}K\left(A+BK\right)^{i}\mathbf{w}_{t-i}+Z_{t}$
@@
The magnitude of this residual term $Z_{t}$ can be bounded, due to the existence of $\frac{\kappa}{\delta}\leq\gamma^{\prime}$ from Exercise 4.6, as
@@
$\leq\sum_{i=h+1}^{t}\kappa(1-\delta)^{i+1}\|K\|\|\mathbf{w}_{t-i}\|$
@@
$=\kappa^{2}\frac{1}{\delta}e^{-\delta h}\leq\varepsilon.$
@@
Similarly, for the corresponding state sequences, first note that
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment. Of particular
Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [Kal60]. The methods for computing such policies are often based on dynamic programming. Often motivated by the need to impose constraints beyond optimality, such as robustness, various other policy parameterizations were introduced. An early example is the Youla parameterization, sometimes called Q-parameterization [YJB76]. Later, the introduction of system-level synthesis (SLS) [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations. For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20, ZZLL23].
The Disturbance Action Control parameterization, introduced in [ABH+19], is similar to system-level synthesis. Like SLS, it is a convex parameterization that is amenable to the tools of convex optimization. However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies. This makes it better suited for machine learning.
The study of regret in the context of control has roots in the work of [AYS11, AYBK14]. These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19, SF20], study regret in a stochastic setting, where the perturbations are stochastic. In such settings, optimal control is well defined. The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19]. The notion of regret as applied to policy classes was implicit in the work of [EDKM04]. The regret metric when applied to policies is different from the counterfactual notion of policy regret defined by [ADT12].
@@ 6.5 Exercises
6.1. Let us say that the total number of time steps $T$ is known in advance. Consider an implementation of online gradient descent where the step size is chosen as $\eta_{t}=D/G\sqrt{T}$, and therefore is uniform over time. Prove that the resultant regret is bounded by $GD\sqrt{T}$, which improves upon Theorem 6.1 by a constant factor. In fact, for ease of exposition, we will use this uniform step size scheme going forward.
6.2. In this exercise, we prove that the optimal solution to a linear control problem can be complex to describe.
@@
    Consider a linear dynamical system in one dimension and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$. Show that the optimal policy in hindsight can have as many domains as loss functions.
@@
    Consider a similar setting in higher dimensions and give a bound on the number of domains as a function of the number of loss functions and the dimension.
6.3. Consider the setting of time-varying linear dynamical systems, i.e.
@@
Prove that any DAC according to definition 6.5 is stable, i.e., prove the analogue of Lemma 6.6 for time-varying systems.
@@
In this chapter, we explore recent advances in machine learning that have significantly reshaped the control problem. These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with, but often outperform traditional approaches. The key characteristics of these modern techniques are as follows:
@@
    Online Learning Over Precomputed Policies Unlike classical control, which relies on precomputing an optimal policy, online control adopts adaptive learning techniques to update policies dynamically. This aligns with the broader framework of adaptive control, where policies evolve in response to observed behavior of the system.
@@
    Improper Learning via Convex Relaxation A fundamental breakthrough in online nonstochastic control is the use of improper learning, where the algorithm competes against a given policy class while learning from a broader class of policies. Using convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, and are accompanied by efficient algorithms with strong theoretical guarantees.
@@ 7.1 From Optimal and Robust to Online Control
Classical control theory has been mainly shaped by two paradigms: optimal control and robust control. These approaches have been instrumental in various engineering applications, but they rely on specific assumptions about disturbances and system dynamics.
@@
Although these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting. Consider the challenge of controlling a drone in an unpredictable environment:
@@
These limitations motivate the need for a more adaptive approach, one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions. This leads us to online nonstochastic control, a fundamentally different paradigm where controllers adapt in real-time without assuming a fixed model for disturbances.
@@ 7.1.1 A Shift to Online Nonstochastic Control
Rather than relying on predefined stochastic or worst-case assumptions, online nonstochastic control introduces a new approach based on sequential decision-making and regret minimization. The core principles are the following.
@@
To build intuition, consider the problem of controlling an autonomous vehicle in an unknown city. If the vehicle had complete foresight of future traffic conditions, it could calculate the optimal route in advance. However, without such knowledge, the vehicle must continuously adjust its route based on real-time observations. Regret measures how far the vehicle’s performance deviates from the best possible path that could have been chosen in hindsight.
@@ 7.2 The Online Nonstochastic Control Problem
In this framework, both disturbances and cost functions are chosen by an adversary, meaning that they may follow no statistically reliable pattern and could even be designed to hinder the controller’s decisions at every step. As a result, the goal is no longer to compute a fixed optimal control policy, but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
A fundamental difficulty in control is the computational tractability of finding the best policy. Classical control methods often rely on dynamic programming, which, in general, suffers from the curse of dimensionality. Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems on trajectories. As pointed out by Rockafellar, finding an optimal control policy in arbitrary convex cost settings can be computationally prohibitive, particularly when state and control constraints are involved.
To address this, online nonstochastic control shifts the focus from computing an optimal policy to competing with a benchmark class of policies using regret minimization. Instead of optimizing a predefined objective function, the controller seeks to minimize its total loss relative to the best policy that could have been chosen in hindsight. This approach bypasses the need to solve an intractable dynamic optimization problem in advance while still achieving performance guarantees over time.
@@
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
@@
7.2.1 Improper Learning and Convex Relaxation
@@
This approach is crucial because enforcing exact constraints on the policy space can lead to computational hardness results, as previously discussed. By relaxing the policy representation, for example, using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
@@ 7.3 The Gradient Perturbation Controller
The algorithm is described in figure (4), for time-varying linear dynamical systems. It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2. The main idea is to learn the parameterization of a DAC policy using known online convex optimization algorithms, namely online gradient descent. We henceforth prove that this parameterization is convex, allowing us to prove the main regret bound.
The notation we use to describe the algorithm is as follows. We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state reached by playing a DAC policy ${M}_{1:h}$ from the beginning of time. Note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$. The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
@@
The GPC algorithm comes with a very strong performance guaranteee: it guarantees vanishing regret vs. the best Disturbance Action Control policy belonging to our reference class in hindsight. According to the relationship between DAC and other policies, this implies vanishing regret versus linear controllers as well. The formal statement is given as follows.
@@
    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable and have a bounded $\ell_{2}$ norm.
@@
According to Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$. However, to directly apply the OGD algorithm to our setting, we need to ensure that the following conditions hold true:
@@
  • We must ensure that optimizing the loss function $\ell_{t}$ as a function of parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
These facts are formalize in the following two lemmas, which we state here and prove at the end of this section.